OpenStudy (anonymous):

Solve for x: please show work :) 8=19x-2x^2

OpenStudy (karatechopper):

Step 1: Factor out the x first Please do so. :)

OpenStudy (anonymous):

can you tell me how to do that?

OpenStudy (anonymous):

\[8=19x-2x^2\] \[2x^2-19x+8=0\] is a start

OpenStudy (karatechopper):

Or yes, can take it Satellite the Legend's way

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

then use the quadratic formula

OpenStudy (anonymous):

I'm not sure what that is

OpenStudy (anonymous):

you cannot factor this using whole numbers

OpenStudy (anonymous):

either of those things

OpenStudy (anonymous):

i've never actually done this before

OpenStudy (karatechopper):

Satellite...wouldn't that just result in a more complex thing? If you factor the x out, or you simply factor the way you said it...it would be more simpler i believe.

OpenStudy (anonymous):

uk

OpenStudy (anonymous):

then your are in a bit of trouble the solutions to \[ax^2+bx+c=0\] are \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok, this is like hieroglyphics to me

OpenStudy (anonymous):

@karatechopper no you can't do that that is why you need the quadratic formula

OpenStudy (anonymous):

ok. i'm lost

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

is there a way you can show me instead? I don't know these terms

OpenStudy (anonymous):

this is your original equation \[8=19x-2x^2\] right?

OpenStudy (karatechopper):

oh wait, i get it. He's right. LIKE ALWAYS.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what now?

OpenStudy (anonymous):

k the first thing we are going to do is set it equal to zero by adding \(2x^2\) to both sides, and subtracting \(19x\) from both sides to start with \[2x^2-19x+8=0\] ok with that step?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i wrote that down. what next?

OpenStudy (anonymous):

the next thing is to realize that this looks just like a general quadratic equation \[ax^2+bx+c=0\] with \[a=2,b=-19,c=8\] is that ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the general solutions to \[ax^2+bx+c=0\] are \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

woe

OpenStudy (anonymous):

that is called the "quadratic formula" and you can't avoid it, you have to use it i will walk you through it

OpenStudy (anonymous):

oh. thanku

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[a=2,b=-19,c=8\] we make a direct substitution and get \[x=\frac{-(-19)\pm\sqrt{(-19)^2-4\times 2\times 8}}{2\times 2}\] let me know when you get what i wrote

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

it is arithmetic from here on in

OpenStudy (anonymous):

i just did it way wrong

OpenStudy (anonymous):

i wrote that down

OpenStudy (anonymous):

and must be missing something

OpenStudy (anonymous):

you get \[x=\frac{19\pm\sqrt{19^2-64}}{4}\] as a start

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

then since \(119^2-64=297\) you get \[x=\frac{19\pm\sqrt{297}}{4}\]

OpenStudy (anonymous):

i don't know what the + sign means when its underlined

OpenStudy (anonymous):

two answers plus or minus

OpenStudy (anonymous):

just like if you have \[x^2=25\] then \(x=\pm5\) it is an abbreviation for \(x=5\) or \(x=-5\)

OpenStudy (anonymous):

ohhhh. that explains alot. ha

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

how dp i know when i'm done?

OpenStudy (anonymous):

when you get \[x=\frac{19\pm\sqrt{297}}{4}\] or if you prefer \[x=\frac{19\pm3\sqrt{33}}{4}\]

OpenStudy (anonymous):

oh, ok. i see. I'm extreme newby

OpenStudy (anonymous):

thank you! that was a great understandable explanation . much appreciated!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Do you know how I can ask a grid question?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

without having a grid to show

OpenStudy (anonymous):

one question is telling u to graph a line, and the other is showing you coordinates on a grid and asking for the slope

OpenStudy (anonymous):

oh, I c attach file